OpenStudy (anonymous):

This table show how long it takes the student in math class to get to school. a).How many students? b).What is the combined time for students to travel to school? c). What is the average time for student to get to school?

OpenStudy (anonymous):

Need the table...

OpenStudy (anonymous):

time 1, 3, 5, 6, 8, 10, 12, 14, 15 student 2, 2, 6, 1, 6, 7, 3 , 2, 1

OpenStudy (anonymous):

^^^^ that is the table..sorry i dont know how to make a table on this

OpenStudy (anonymous):

total student = 2+2+6+1+6+7+3+2+2 = answer for a total time = (1+3+5+6+8+10+12+14+15)/9 = answer for c